A person runs with an initial velocity of 5.0 m/s. Then he slows down at a constant rate of -1.2 m/s2 over a period of 3 seconds. What is his velocity at the end of time.


1
Expert's answer
2021-10-18T17:41:29-0400

"a=\frac{v_2-v_1}{t}=1.2" m/s2


"v_2=v_1+at=5-1.2\cdot 3=1.4" m/s